# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

### v1

@@ Abstract
In this paper, we explore the concept of integrated sensing and communication (ISAC) within a downlink cell-free massive MIMO (multiple-input multiple-output) system featuring multi-static sensing and users requiring ultra-reliable low-latency communications (URLLC). Our focus involves the formulation of two non-convex algorithms that jointly solve power and blocklength allocation for end-to-end (E2E) minimization. The objectives are to jointly minimize sensing/communication processing and transmission energy consumption, while simultaneously meeting the requirements for sensing and URLLC. To address the inherent non-convexity of these optimization problems, we utilize techniques such as the Feasible Point Pursuit - Successive Convex Approximation (FPP-SCA), Concave-Convex Programming (CCP), and fractional programming. We conduct a comparative analysis of the performance of these algorithms in ISAC scenarios and against a URLLC-only scenario where sensing is not integrated. Our numerical results highlight the superior performance of the E2E energy minimization algorithm, especially in scenarios without sensing capability. Additionally, our study underscores the increasing prominence of energy consumption associated with sensing processing tasks as the number of sensing receive access points rises. Furthermore, the results emphasize that a higher sensing signal-to-interference-plus-noise ratio threshold is associated with an escalation in E2E energy consumption, thereby narrowing the performance gap between the two proposed algorithms.
@@ I Introduction
Integrated sensing and communication (ISAC) aims to enhance spectral and hardware efficiency, offering potential applications in 6G communication networks, including autonomous vehicles, smart factories, and location/environment-aware scenarios [1, 2]. Particularly, in ultra-reliable low-latency communication (URLLC) use cases like industrial Internet-of-Things (IIoT) and vehicle-to-everything (V2X) communications, ISAC can play a crucial role. For instance, in V2X networks, ISAC systems can detect targets such as pedestrians and vehicles, delivering sensing information reliably and with minimal delay for URLLC users like autonomous vehicles [3]. Thus, a comprehensive study of these technologies together is imperative to enable mission-critical applications in future wireless networks.
For many URLLC applications, such as V2X and IIoT, short codewords are needed to satisfy latency constraints where codes with short blocklengths, e.g., 100 symbols are employed. Therefore, the finite blocklength regime needs to be considered to model decoding error probability (DEP) [4, 5, 6]. In such applications, short data packets should be delivered with a minimum reliability of 99.999% and an end-to-end (E2E) delay of less than 10-150 ms [7, 8]. Therefore, optimizing the blocklength can play an important role to meet the reliability and delay requirements.
Conversely, while ISAC introduces perceptive mobile networks through network-level sensing [9, 10, 11, 12], it is anticipated that ISAC networks may consume more energy compared to communication networks due to sensing tasks. This motivates a comprehensive study of the end-to-end (E2E) energy consumption of ISAC networks taking into account not only energy consumption for transmission but also energy consumption for processing tasks. Centralized radio access network (C-RAN) architecture has been already introduced as an energy-efficient solution for the mobile networks. There is also an inherent connection between cell-free massive MIMO (multiple-input multiple-output) and the C-RAN architecture due to the joint transmission and processing the UEs’ signals[13]. However, E2E energy-aware resource orchestration is a critical aspect that must be considered when optimizing cell-free massive MIMO systems on C-RAN architectures to fully benefit from energy savings potential [14].
I-A Related Work
@@
To the best of our knowledge, there are few works that jointly consider URLLC and ISAC. In [3], a joint precoding scheme is proposed to minimize transmit power, satisfying sensing and delay requirements. Moreover, joint ISAC beamforming and scheduling design is addressed in [15] for periodic and aperiodic traffic. The existing works consider a single ISAC base station, while cell-free massive MIMO with multiple distributed access points (APs) has been shown to facilitate URLLC [6, 16, 17] and ISAC[18, 19] implementations.
@@
In the conference version [20] of this work, we studied for the first time ISAC-enabled cell-free massive MIMO with URLLC users. In this paper, we extend [20] from an E2E energy minimization perspective. The consideration of E2E energy-awareness has been explored in various contexts, as reflected in prior works such as [21, 22, 23, 14, 24]. In particular, [14] studied fully virtualized E2E power minimization problem for cell-free massive MIMO on open radio access network (O-RAN) architecture by taking the radio, fronthaul, and processing resources into account. However, this work neither considered ISAC nor URLLC communications.
@@ I-B Contributions
In this paper, we study E2E energy consumption in a cell-free massive MIMO system with URLLC user equipments (UEs) and multi-static sensing in a cluttered environment. Both sensing and communication signals contribute to the sensing task. However, the sensing signals can cause interference for the UEs. Thus, we design the precoding vectors to null the interference only for the UEs. We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements and a minimum signal-to-interference-plus-noise ratio (SINR) as the sensing requirement. The latter is motivated by the fact that given a fixed false alarm probability, detection probability increases with a higher sensing SINR [25], [26, Chap. 3 and 15]. In the conference version of this work [20], we propose a successive convex approximation-based power allocation algorithm that maximizes energy efficiency while satisfying the sensing and URLLC requirements for a given blocklength. However, the impact of blocklength optimization is not considered in [20].
Different from the conference paper, we conduct a comprehensive analysis of the processing requirements, specifically in terms of giga operations per second (GOPS), and examine its impact on processing energy consumption. The interplay between processing requirements and blocklength yields intriguing results, as demonstrated by the simulation findings presented in this paper.
The main contributions of this paper are outlined as follows:
  • We derive an upper bound for the DEP and transmission delay in finite blocklength cell-free massive MIMO communications.
  • We propose two joint blocklength and power allocation algorithms, which minimize the transmission energy consumption and E2E energy consumption, respectively while meeting URLLC and sensing requirements. The former algorithm is represented by SeURLLC+ and the latter is represented by E2E URLLC+. To solve the non-convex optimization problems, we benefit from the Feasible Point Pursuit - Successive Convex Approximation (FPP-SCA), Concave-Convex Programming (CCP), and fractional programming.
  • We investigate the impact of E2E energy minimization, encompassing both radio and processing energy consumption, in the above-mentioned ISAC scenario, as well as a scenario where only URLLC requirements are taken into account. The algorithms for URLLC-only scenarios are represented by URLLC-only and E2E URLLC-only. The former corresponds to transmission energy minimization and the latter corresponds to E2E energy minimization.
  • We conduct a sensitivity analysis by investigating the impact of DEP threhsold, maximum delay threshold, and sensing SINR threshold on E2E energy consumption.
The rest of the paper is organized as follows. In Section II, we present the system model. In Section III, we provide URLLC analysis accounting for decoding error probability and delay in finite blocklength regime. Section IV is devoted to the sensing analysis. In Section V, we describe the E2E power model and derive the GOPS analysis for both communication and sensing. Next, we present our optimization problems in Section VI. Finally, the numerical results and conclusion are provided in Section VII and Section VIII, respectively.
@@ II System Model
We consider an ISAC-enabled cell-free massive MIMO system in URLLC scenarios. The system adopts a centralized radio access network (C-RAN) architecture [27] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig. 1. All the APs are interconnected via fronthaul links to the edge cloud and operate in full synchronization. We consider the original form of cell-free massive MIMO [6], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
Figure 1: Illustration of the ISAC system in cell-free massive MIMO with URLLC UEs.
@@
We consider the finite blocklength regime for URLLC UEs, where a packet of $b_{i}$ bits is sent to UE $i$ within a transmission block with blocklength $L=L_{p}+L_{d}$ symbols using the coherence bandwidth $B$. $L_{p}$ and $L_{d}$ are the number of symbols for pilot and data, respectively. It is expected that duration of each URLLC transmission, denoted by $T$, is shorter than one coherence time $T_{c}$, i.e., $T<T_{c}$ [17]. Without loss of generality, we assume the coherence time is equal to URLLC transmission time so that in for each transmission we first estimate the channel.
II-A Channel Modeling and Downlink Communications
@@
$\textbf{h}_{i}=\begin{bmatrix}\textbf{h}_{i,1}^{T}&\ldots&\textbf{h}_{i,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ (3)
for UE $i$.
$\vspace{-1mm}\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\boldsymbol{\rho},\vspace{-1mm}$ (4)
where the vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for transmit AP $k$ corresponding to UE $i$ and the sensing signal, respectively. In (4), $\textbf{W}_{k}=\begin{bmatrix}\textbf{w}_{0,k}&\textbf{w}_{1,k}&\cdots&\textbf{w}_{N_{\rm ue},k}\end{bmatrix}$, ${\textbf{D}}_{\rm s}[m]=\textrm{diag}\left(s_{0}[m],s_{1}[m],\ldots,s_{N_{\rm ue}}[m]\right)$ is the diagonal matrix containing the sensing and communication symbols, and $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$.
Moreover, let $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ be the independent receiver noise at UE $i$ at time instance $m$. The received signal at UE $i$ is given as
@@
where the collective precoding vectors
@@
are the centralized precoding vectors. The vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for ISAC AP $k$ corresponding to UE $i$ and the sensing signal, respectively.
@@ II-B LMMSE-Based Channel Estimation
Within the URLLC transmission time comprising $L$ symbols, $L_{p}<L$ symbols are allocated for pilot transmission. A collection of $L_{p}$ mutually orthogonal pilot sequences is represented by $\boldsymbol{\phi}_{1},\ldots,\boldsymbol{\phi}_{L_{p}}\in\mathbb{C}^{L_{p}}$, where $\|\boldsymbol{\phi}_{t}\|^{2}=L_{p}$ for $t=1,\ldots,L_{p}$. These pilot sequences are assigned to the $N_{\rm ue}$ UEs. If the number of pilot sequences is less than the number of UEs (i.e., $L_{p}<N_{\rm ue}$), then each pilot sequence may be assigned to multiple UEs. Let $t_{i}\in\{1,\ldots,L_{p}\}$ denote the index of the pilot sequence that is assigned to UE $i$ and we define $\mathcal{P}_{i}=\{j\in\{1,\ldots,N_{\rm ue}\}|t_{j}=t_{i}\}$ as the set of UEs that share the same pilot with UE $i$.
@@
$\textbf{Y}_{k}^{p}=\sum_{i=1}^{N_{\rm ue}}\sqrt{\eta_{i}}\textbf{h}_{i,k}\boldsymbol{\phi}_{t_{i}}^{T}+\textbf{N}_{k},$ (9)
where $\eta_{i}>0$ is the transmit power by UE $i$ and $\textbf{N}_{k}\in\mathbb{C}^{M\times L_{p}}$ is the received noise at AP $k$ with independent and identically distributed (i.i.d.) $\mathcal{CN}(0,\sigma_{n}^{2})$ entries. We assume that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$, representing the long-term statistics for the channel in (1), are known while the random phase-shift $\varphi_{i,k}$ is unknown. With respect to estimating $\textbf{h}_{i,k}$, let us form $\textbf{y}_{t_{i},k}^{p}=\frac{1}{\sqrt{L_{p}}}\textbf{Y}_{k}^{p}\boldsymbol{\phi}_{t_{i}}^{*}\in\mathbb{C}^{M}$, where
@@
where $\textbf{n}_{t_{i},k}=\frac{1}{\sqrt{L_{p}}}\textbf{N}_{k}\boldsymbol{\phi}_{t_{i}}^{*}\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}\textbf{I}_{M})$. If the number of UEs is greater than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [27, Alg. 4.1]. The LMMSE estimate of $\textbf{h}_{i,k}$ is given by [29]
$\hat{\textbf{h}}_{i,k}=\sqrt{\eta_{i}L_{p}}\textbf{R}_{i,k}^{\prime}\left(\boldsymbol{\Psi}^{\prime}_{t_{i},k}\right)^{-1}\textbf{y}_{t_{i},k}^{p},$ (11)
where $\textbf{R}_{i,k}^{\prime}$ and $\boldsymbol{\Psi}_{t_{i},k}^{\prime}$ are the spatial correlation matrix corresponding to channel $\textbf{h}_{i,k}$ and the correlation matrix of the received signal, respectively, given by
@@
$\boldsymbol{\Psi}^{\prime}_{t_{i},k}\triangleq\sum_{j\in\mathcal{P}_{i}}\eta_{j}L_{p}\textbf{R}_{j,k}^{\prime}+\sigma_{n}^{2}\textbf{I}_{M}.$ (13)

### v2

@@ Abstract
In this paper, we explore the concept of integrated sensing and communication (ISAC) within a downlink cell-free massive MIMO (multiple-input multiple-output) system featuring multi-static sensing and users requiring ultra-reliable low-latency communications (URLLC). Our work focuses on devising two algorithms that concurrently optimize power and blocklength allocation. The objectives of these algorithms are to minimize the energy consumption solely for transmission and the overall end-to-end (E2E) energy consumption, which encompasses sensing/communication processing and transmission. This is achieved while meeting the necessary requirements for sensing and URLLC. To address the inherent non-convexity of these optimization problems, we apply feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming. Our numerical analysis compares the performance of these algorithms in ISAC scenarios and against a baseline URLLC-only scenario without sensing integration. Results show the E2E energy minimization algorithm’s effectiveness, particularly in scenarios without sensing. Additionally, our study underscores the increasing prominence of energy consumption associated with sensing processing tasks as the number of sensing receive access points rises, indicating potential for energy savings by selectively deactivating APs. Furthermore, the results emphasize that although increasing the number of transmit APs can improve URLLC satisfaction and sensing update rate by reducing blocklength, careful management is essential to avoid excessive energy consumption and reduced detection probability.
@@ I Introduction
6G mobile networks are expected to offer various sensing-based applications such as autonomous vehicles, smart homes/cities/factories, remote healthcare, and location/environment-aware scenarios [1, 2]. Many of these applications, such as autonomous vehicles and industrial Internet-of-things (IIoT), demand ultra-reliable delivering of the sensing information to the user equipments (UEs) with a minimum reliability of 99.999% and within a low end-to-end (E2E) delay of less than 10-150 ms [3, 4, 5]. This demand necessitates two key aspects defined in 6G networks: integrated sensing and communication (ISAC) together with ultra-reliable low-latency communication (URLLC). For most URLLC applications, short codewords are usually needed to satisfy latency constraints where codes with short blocklengths, e.g., 50-400 symbols are employed. Short blocklengths result in decoding errors; therefore, to ensure reliability, it is essential to consider the finite blocklength regime for modeling the decoding error probability (DEP) [6, 7, 8].
In mission-critical applications such as traffic control and autonomous vehicles, real-time transmission of sensing information to URLLC UEs is essential, as the speed of information updates directly impacts the effectiveness of these applications. In this paper, we introduce the concept of the “sensing update rate (SUR),” which quantifies the number of sensing measurements and updates per time unit [9]. The SUR is influenced by the blocklength, with longer blocklengths resulting in lower update rates due to the extended duration of each sensing operation. This dependency creates an opportunity to optimize blocklength to balance reliable communication with high SUR. Although URLLC and ISAC have typically been studied as separate domains, a joint design and optimization approach is required to meet the distinct requirements of both and maximize integration gains.
Cell-free massive multiple-input multiple-output (CF-mMIMO) has been shown to be a good candidate to facilitate stringent requirements of URLLC applications [8, 10, 11] as it provides high reliability by handling large path-loss variations and improves cell-edge user performance. CF-mMIMO can also facilitate ISAC implementations leveraging distributed access points (APs) to deploy bi-static/ multi-static sensing eliminating the need for full-duplex capability at the APs [12, 13, 14, 15]. The central processing of network functions within a centralized radio access network (C-RAN) architecture facilitates coordination among distributed APs, enabling joint transmission and processing in CF-mMIMO[16]. C-RAN has evolved into virtualized C-RAN and, with the open interfaces, further into Open RAN (O-RAN), where virtualization enables the use of general-purpose servers for jointly processing of baseband and sensing signals.
However, deploying these technologies in 6G is anticipated to result in substantial energy consumption. In CF-mMIMO networks, baseband processing combined with sensing tasks is expected to lead to higher energy usage in ISAC networks compared to conventional communication systems. The impact of sensing on energy consumption in such networks has not been well-studied. Existing studies on green CF-mMIMO primarily focus on optimizing transmit power consumption in radio access, with limited work addressing processing energy, aside from [17]. Virtualized C-RAN architecture facilitates centralized processing and enables the sharing of processing resources between APs and tasks, paving the way for end-to-end (E2E) optimization of both processing and radio resources in ISAC networks. This motivates a comprehensive study of the E2E energy consumption in ISAC networks, considering not only energy consumption for transmission but also for processing tasks. The primary research question addressed in this paper is how to incorporate sensing in CF-mMIMO systems with URLLC while minimizing E2E energy consumption.
I-A Literature Review
The finite blocklength regime and URLLC have primarily been studied in the context of cellular networks (see [18, 19, 20] and references therein), with limited research addressing URLLC in CF-mMIMO networks. The potentials of massive MIMO in meeting URLLC requirements for industrial automation are investigated in [21, 22, 23]. Network availability is investigated in [21] and [22], and [23] proposes a power allocation algorithm to maximize the minimum signal-to-interference-plus-noise ratio (SINR). The authors in [24] propose two power optimization algorithms to provide URLLC for traditional ground UEs and unmanned aerial vehicles (UAVs). The objectives in [24] are maximizing either sum or minimum URLLC rate and they study neither energy minimization nor blocklength optimization. Max-min rate optimization is also studied in [11], where the authors adapt a special class of the conjugate beamforming for a CF-mMIMO with single-antenna APs. Moreover, the authors aim to address an energy efficiency optimization problem.
Joint power and blocklength optimization is studied in [25] and [26] to optimize the reliability in a factory automation scenario and heterogeneous cellular networks, respectively. Cell-free networks as well as energy minimization are not studied in these works. Moreover, sensing requirements are not considered in [23, 21, 24, 11, 25, 26].
@@ I-A Literature Review
There are few works that jointly consider URLLC and ISAC. In [5], a joint precoding scheme is proposed to minimize transmit power, satisfying sensing and delay requirements. Moreover, joint ISAC beamforming and scheduling design is addressed in [27] for periodic and aperiodic traffic.
@@
The consideration of E2E energy-awareness has been explored in various contexts, as reflected in prior works such as [28, 29, 17, 30, 31]. In particular, [30] studied fully virtualized E2E power minimization problem for CF-mMIMO on O-RAN architecture by taking the radio, fronthaul, and processing resources into account. Joint user scheduling and power allocation scheme for cell-free ISAC systems is studied in [32] that aims to maximize the sum-rate of the communication UEs and the sensing targets. However, the authors do not consider either URLLC requirements or energy minimization.
To the best of our knowledge, for the first time, ISAC in CF-mMIMO with URLLC UEs and with the focus on energy efficiency was studied in our previous work[33]. In [33], we propose a successive convex approximation-based power allocation algorithm that maximizes energy efficiency while satisfying the sensing and URLLC requirements for a given blocklength. However, the impact of blocklength optimization is not considered in [33]. In this paper, we extend [33] from an E2E energy minimization perspective and aim to address a joint blocklength and power optimization problem. Here, E2E refers to considering energy consumption at both radio site and the central cloud.
@@ I-B Contributions
In this paper, we study E2E energy consumption in a CF-mMIMO system with URLLC UEs and multi-static sensing in a cluttered environment. We sense one candidate location of the target utilizing communication signals and an additional sensing signal. The sensing signals can cause interference for the UEs. Thus, we design the sensing precoding vectors to null the interference for the UEs. We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements and a minimum SINR as the sensing requirement. The latter is motivated by the fact that given a fixed false alarm probability, detection probability increases with a higher sensing SINR [34], [9, Chap. 3 and 15].
Different from [33], we conduct a comprehensive analysis of the processing requirements, specifically in terms of giga-operations per second (GOPS), and examine its impact on processing energy consumption. The main contributions of this paper are outlined as follows:
  • We propose two joint blocklength and power allocation algorithms aimed at minimizing transmission energy consumption and E2E energy consumption, respectively, while satisfying URLLC and sensing requirements. To address the non-convex optimization challenges, we employ feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming.
  • We analyze the impact of E2E energy minimization, covering both radio and processing energy consumption, within the proposed ISAC scenario and a comparative scenario that considers only URLLC requirements.
  • We examine two distinct target detection schemes, comparing their performance in terms of E2E energy consumption, detection probability, and sensing update rate.
  • We conduct sensitivity analyses to evaluate the effects of DEP threshold, sensing SINR threshold, and the number of ISAC transmit APs and sensing receive APs on E2E energy consumption, detection probability, and sensing update rate.
The rest of the paper is organized as follows: Section II introduces the system model. Section III provides URLLC analysis, considering DEP and delay in finite blocklength regime. Section IV covers the sensing analysis. Section V describes the E2E power model and derives GOPS analysis for both communication and sensing. Optimization problems are presented in Section VI, followed by numerical results and conclusions in Sections VII and VIII, respectively.
@@ II System Model
We study ISAC in a CF-mMIMO system in URLLC scenarios. The system adopts a virtualized C-RAN architecture [35] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig. 1. All the APs are interconnected via fronthaul links to the edge cloud and operate in full synchronization. We consider the original form of CF-mMIMO [8], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
Figure 1: Illustration of an ISAC system in CF-mMIMO with URLLC UEs.
@@
We consider the finite blocklength regime for URLLC UEs, where a packet of $b_{i}$ bits is sent to UE $i$ within a transmission block with blocklength $L=L_{p}+L_{d}$ symbols using the coherence bandwidth $B$. $L_{p}$ and $L_{d}$ are the number of symbols for pilot and data, respectively. It is expected that duration of each URLLC transmission, denoted by $T$, is shorter than one coherence time $T_{c}$, i.e., $T<T_{c}$ [11]. Without loss of generality, we assume that we have a channel estimation phase in each transmission.
II-A Downlink Communication
$\vspace{-1mm}\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\bm{\rho},\vspace{-1mm}$ (1)
where the vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for transmit AP $k$ corresponding to UE $i$ and the sensing signal, respectively. In (1), $\textbf{W}_{k}=\begin{bmatrix}\textbf{w}_{0,k}&\textbf{w}_{1,k}&\cdots&\textbf{w}_{N_{\rm ue},k}\end{bmatrix}$, ${\textbf{D}}_{\rm s}[m]=\textrm{diag}\left(s_{0}[m],s_{1}[m],\ldots,s_{N_{\rm ue}}[m]\right)$ is the diagonal matrix containing the sensing and communication symbols, and $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$.
@@ II-A Downlink Communication
$\textbf{h}_{i}=\begin{bmatrix}\textbf{h}_{i,1}^{T}&\ldots&\textbf{h}_{i,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M},$ (3)
for UE $i$. The received signal at UE $i$ is given as
@@
where $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ is the independent receiver noise at UE $i$ at time instance $m$ and the collective precoding vectors
@@
are the centralized precoding vectors.
@@ II-B LMMSE-Based Channel Estimation
Within the URLLC transmission time comprising $L$ symbols, $L_{p}<L$ symbols are allocated for pilot transmission. A collection of $L_{p}$ mutually orthogonal pilot sequences is represented by $\bm{\phi}_{1},\ldots,\bm{\phi}_{L_{p}}\in\mathbb{C}^{L_{p}}$, where $\|\bm{\phi}_{t}\|^{2}=L_{p}$ for $t=1,\ldots,L_{p}$. These pilot sequences are assigned to the $N_{\rm ue}$ UEs. If the number of pilot sequences is less than the number of UEs (i.e., $L_{p}<N_{\rm ue}$), then each pilot sequence may be assigned to multiple UEs. Let $t_{i}\in\{1,\ldots,L_{p}\}$ denote the index of the pilot sequence that is assigned to UE $i$ and we define $\mathcal{P}_{i}=\{j\in\{1,\ldots,N_{\rm ue}\}|t_{j}=t_{i}\}$ as the set of UEs that share the same pilot with UE $i$.
@@
$\textbf{Y}_{k}^{p}=\sum_{i=1}^{N_{\rm ue}}\sqrt{\eta_{i}}\textbf{h}_{i,k}\bm{\phi}_{t_{i}}^{T}+\textbf{N}_{k},$ (8)
where $\eta_{i}>0$ is the transmit power by UE $i$ and $\textbf{N}_{k}\in\mathbb{C}^{M\times L_{p}}$ is the received noise at AP $k$ with independent and identically distributed (i.i.d.) $\mathcal{CN}(0,\sigma_{n}^{2})$ entries. We assume that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$, representing the long-term statistics for the channel in (2), are known while the random phase-shift $\varphi_{i,k}$ is unknown. With respect to estimating $\textbf{h}_{i,k}$, let us form $\textbf{y}_{t_{i},k}^{p}=\textbf{Y}_{k}^{p}\bm{\phi}_{t_{i}}^{*}\in\mathbb{C}^{M}$, where
@@
where $\textbf{n}_{t_{i},k}=\textbf{N}_{k}\bm{\phi}_{t_{i}}^{*}\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}L_{p}\textbf{I}_{M})$. If the number of UEs is greater than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [35, Alg. 4.1]. The LMMSE estimate of $\textbf{h}_{i,k}$ is given by [37]
$\hat{\textbf{h}}_{i,k}=\sqrt{\eta_{i}}\textbf{R}_{i,k}^{\prime}\left(\bm{\Psi}^{\prime}_{t_{i},k}\right)^{-1}\textbf{y}_{t_{i},k}^{p},$ (10)
where $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$ are the spatial correlation matrix corresponding to channel $\textbf{h}_{i,k}$ and the correlation matrix of the received signal, respectively, given by
@@
$\bm{\Psi}^{\prime}_{t_{i},k}\triangleq\sum_{j\in\mathcal{P}_{i}}\eta_{j}L_{p}\textbf{R}_{j,k}^{\prime}+\sigma_{n}^{2}\textbf{I}_{M}.$ (12)